Triangle ABC is a right triangle. If side AC=5 and side BC=6, What is the measure of side AB?
Answer by checkley71(8403)

IF AB IS THE HYPOTENUSE THEN
AB^2=5^2+6^2
AB^2=25+36
AB^2=61
AB=SQRT61
AB=7.81
---------------------
IF BC IS THE HYPOTENUSE THEN
6^2=5^2+AB^2
36=25+AB^2
36-25=AB^2
9=AB^2
AB=SQRT9
AB=3
------------------------------